# On the Validity of the Background Subtraction Method for Black Hole Thermodynamics in Matter–Coupled Gravity Theories

### v2

@@ II The derivation of background subtraction method from Iyer–Wald formalism
$ds^{2}=\!-\!\Big(1\!-\!\frac{2m}{r}\!-\!\frac{\Lambda r^{2}}{3}\Big)dt^{2}+\frac{dr^{2}}{1\!-\!\frac{2m}{r}\!-\!\frac{\Lambda r^{2}}{3}}+r^{2}d\Omega_{2}^{2}.$ (13)
@@
$\bm{\omega}\equiv\delta\mathbf{Q}_{\xi_{t}}-\xi_{t}\cdot\bm{\Theta}[\delta\phi].$ (15)
Integrating $d\bm{\omega}$ over a spatial hypersurface $\Sigma$ that extends from the bifurcation surface $S_{r_{h}}$ to a large-$r$ surface $S_{\infty}$, Gauss’s law yields
$\int_{S_{r_{h}}}\bm{\omega}=\int_{S_{\infty}}\bm{\omega}.$ (16)
@@
The left-hand side of Eq. (16) is known to equal $T\,\delta S$, where $T$ is the Hawking temperature and $S$ is the black hole entropy Wald:1993nt ; Iyer:1994ys . For the right-hand side, we can evaluate $\bm{\Theta}$ on the timelike boundary $\Gamma$ at infinity, where $S_{\infty}=\Sigma\cap\Gamma$ and the induced volume form $\hat{\bm{\epsilon}}$ on $\Gamma$ satisfies $\bm{\epsilon}=\bm{n}\wedge\hat{\bm{\epsilon}}$ (where $\bm{n}\propto\nabla r$ being the outward unit normal vector to $\Gamma$). Following standard procedures Padmanabhan:2014lwa ; Chakraborty:2017zep ; Jiang:2018sqj , $\bm{\Theta}$ at the boundary can be decomposed as
@@
Substituting $\int_{S_{h}}\bm{\omega}=T\delta S$ and Eqs. (15), (17) into Eq. (16), we obtain
@@
$ds^{2}=-\left(1-\frac{\Lambda r^{2}}{3}\right)\lambda(r_{c})^{2}dt^{2}+\frac{dr^{2}}{1-\frac{\Lambda r^{2}}{3}}+r^{2}d\Omega_{2}^{2}.$ (21)
To ensure the black hole spacetime and the background behave consistently at $r=r_{c}\rightarrow\infty$, the redshift factor $\lambda(r_{c})$ should be chosen as the form
@@
For this horizonless background spacetime, we have $0=\int_{S_{\infty}}\bm{\omega}$, leading to
@@
For the Schwarzschild-AdS black hole example, one can insert the expressions of $\bm{Q}_{\xi_{t}}$ and $\bm{B}$ into Eq. (26) and verify that it indeed gives the expected mass $M=m$. This expression can also be rewritten in the ADM or Brown–York form, but that lies beyond our present scope; see Iyer:1994ys ; Harlow:2019yfa ; Guo:2024oey for details.
@@
All of the above applies to both asymptotically AdS and asymptotically flat spacetimes. In the latter case, the boundary term $\int_{\partial\mathcal{M}}^{(\mathrm{Reg.})}\bm{B}_{E}$ provides the dominant contribution to $I_{E}$. However, for asymptotically AdS black holes in Einstein gravity and higher-derivative gravity theories, Ref. Guo:2025ohn shows that the boundary term always vanishes after subtraction, following a careful examination of the asymptotic behaviors of such terms. This means that when using Eqs. (26) and (28) to compute the mass and Euclidean action, one may safely ignore $\bm{B}_{E}$. We remark that this property holds for pure gravity in AdS spacetime. Once we go beyond this regime, such as in matter-coupled theories, we should carefully re-examine possible boundary contributions.
@@ III Applying the subtraction method to matter–coupled gravity
In last section, we have shown that for pure gravity, the background subtraction method directly follows from the Iyer–Wald formalism and yields correct thermodynamics. The reasoning is quite general, so there is no obvious reason why this conclusion should fail when matter fields are present. For a diffeomorphism-invariant Lagrangian, the inclusion of matter only modifies the explicit expressions of $\bm{Q}$ and $\bm{\Theta}$ by adding matter-related contributions, and none of the general steps in Sec. II are altered.
@@ III.1 Higher-derivative gravity with non-minimally coupled Maxwell field
$\bm{\omega}=\delta\mathbf{Q}_{\xi}-\xi\cdot\bm{\Theta}[\delta\phi],$ (38)
@@
We now take the timelike Killing vector $\xi_{t}=\frac{\partial}{\partial t}$. The identity $\int_{S_{h}}\bm{\omega}=\int_{S_{\infty}}\bm{\omega}$ produces
@@
Next, we introduce the reference background spacetime. For the metric in Eq. (34), we set $r_{h}=q=0$ and apply a redshift factor to ensure the black hole spacetime and background behave consistently as $r=r_{c}\to\infty$. For this horizonless background, the identity $\int_{S_{\infty}}\bm{\omega}=0$ holds, leading to:
@@ III.3 Two subtleties in applying the Iyer–Wald formalism and background subtraction method
In matter-coupled gravity, variations of matter field parameters contribute to $\bm{\omega}=\delta\mathbf{Q}_{\xi_{t}}-\xi_{t}\cdot\bm{\Theta}[\delta\phi]$, which can lead to non-integrable variations, i.e., expressions that cannot be written as the differential of a single quantity.
For gauge fields, this issue is often resolved by gauge choice: for example, if the gauge field $A_{\mu}$ is required to vanish at the horizon, one finds $\int_{\infty}\bm{\omega}=\delta M-\Phi\delta Q$, which is a non-integrable variation. But redefining the gauge can shift the $\Phi\delta Q$ term from the asymptotic boundary to the horizon, neatly yielding $\int_{\infty}\bm{\omega}=\delta M$.
@@
$\int_{\infty}\bm{\omega}\sim-2\delta f_{v}+\phi_{s}\delta\phi_{v}+2\phi_{v}\delta\phi_{s},$ (71)
@@
This non-integrability does not indicate a failure of the Iyer–Wald formalism and the subtraction method. Actually, they still yield the correct thermodynamic relations among variations of different quantities involving $\delta f_{v}$, $\delta\phi_{v}$, and $\delta\phi_{s}$. The main problem is that ambiguities arise when identifying $\delta M$ from $\int_{\infty}\bm{\omega}$: neither the Iyer–Wald formalism nor the background subtraction method provides a clear criterion for classifying these terms. Accordingly, one must resort to an alternative approach. In this regard, Ref. Li:2020spf resolves the issue using holographic renormalization, properly rewriting the integral (71) as
$\int_{\infty}\bm{\omega}\sim\delta\mathcal{E}+\langle O\rangle d\phi_{s},$ (74)
@@ III.3.2 Discrepancy between black hole entropy and Wald entropy
The black hole entropy $S$ is defined via $\int_{S_{h}}\bm{\omega}=T\delta S$, and the Wald entropy $S_{W}$ is defined via $\int_{S_{h}}\mathbf{Q}_{\xi}=TS_{W}$. These two quantities coincide with each other, differing only in extreme cases where matter fields diverge at the horizon An:2024fzf ; Feng:2015oea .
Concretely, the equality between $S$ and $S_{W}$ relies on the horizon Killing vector vanishing at the bifurcation surface, which eliminates the $\xi\cdot\bm{\Theta}[\delta\phi]$ term in $\bm{\omega}$. However, if a matter field diverges at $r=r_{h}$, it can counteract this effect and yield a non-trivial contribution to $\int_{S_{h}}\bm{\omega}$, causing the inequality $S\neq S_{W}$. When this situation arises, it brings about difficulties in consistently interpreting black hole thermodynamics. The awkward problem of $S_{W}\neq S$ (within the Iyer–Wald formalism) translates to $I_{E}=\beta(M-TS_{W})\neq\beta(M-TS)$ (within the Euclidean action method). In such cases, $I_{E}$ cannot be legitimately related to the system’s free energy, and we thus cannot extract the various thermodynamic quantities of the black hole from $I_{E}$.
@@ Acknowledgments
YX would like to thank Hong-Bao Zhang, Xi-Yao Guo and Jie-Qiang Wu for useful discussions. YX is also thankful to the Higgs Centre for Theoretical Physics at the University of Edinburgh for providing research facilities and hospitality during the visit. This work was supported in part by China Scholarship Council with Grant No. 202408130101, the Hebei Natural Science Foundation with Grant No. A2024201012, and the National Natural Science Foundation of China with Grant No. 12475048.

### v3

# Validity of the Background Subtraction Method for Black Hole Thermodynamics in Matter–Coupled Gravity Theories

@@ II The derivation of background subtraction method from Iyer–Wald formalism
$ds^{2}=\!-\!\Big(1\!-\!\frac{2m}{r}\!-\!\frac{\Lambda r^{2}}{3}\Big)dt^{2}+\frac{dr^{2}}{1\!-\!\frac{2m}{r}\!-\!\frac{\Lambda r^{2}}{3}}+r^{2}d\Omega_{2}.$ (13)
@@
$\bm{k}_{\xi_{t}}\equiv\delta\mathbf{Q}_{\xi_{t}}-\xi_{t}\cdot\bm{\Theta}[\delta\phi].$ (15)
Integrating $d\bm{k}_{\xi_{t}}$ over a spatial hypersurface $\Sigma$ that extends from the bifurcation surface $S_{r_{h}}$ to a large-$r$ surface $S_{\infty}$, Gauss’s law yields
$\int_{S_{r_{h}}}\bm{k}_{\xi_{t}}=\int_{S_{\infty}}\bm{k}_{\xi_{t}}.$ (16)
@@
For the static case where $\xi_{t}$ is the horizon Killing vector, the left-hand side of Eq. (16) is known to equal $T\,\delta S$, where $T$ is the Hawking temperature and $S$ is the black hole entropy Wald:1993nt ; Iyer:1994ys . For the right-hand side, we can evaluate $\bm{\Theta}$ on the timelike boundary $\Gamma$ at infinity, where $S_{\infty}=\Sigma\cap\Gamma$ and the induced volume form $\hat{\bm{\epsilon}}$ on $\Gamma$ satisfies $\bm{\epsilon}=\bm{n}\wedge\hat{\bm{\epsilon}}$ (where $\bm{n}\propto\nabla r$ being the outward unit normal vector to $\Gamma$). Following standard procedures Padmanabhan:2014lwa ; Chakraborty:2017zep ; Jiang:2018sqj , $\bm{\Theta}$ at the boundary can be decomposed as
@@
Substituting $\int_{S_{h}}\bm{k}_{\xi_{t}}=T\delta S$ and Eqs. (15), (17) into Eq. (16), we obtain
@@
$ds^{2}=-\left(1-\frac{\Lambda r^{2}}{3}\right)\lambda(r_{c})^{2}dt^{2}+\frac{dr^{2}}{1-\frac{\Lambda r^{2}}{3}}+r^{2}d\Omega_{2}.$ (21)
At $r=r_{c}\rightarrow\infty$, the redshift factor $\lambda(r_{c})$ should be chosen as the form
@@
The redshift factor (22) is widely used in the literature for evaluating the Euclidean action Gibbons:1976ue ; hawkingpage ; Gubser:1998nz ; Gibbons:2004ai ; Dutta:2006vs ; Xiao:2023two , with the key physical motivation for this choice being that it ensures the black hole spacetime and pure AdS spacetime have identical asymptotic geometries (i.e., identical induced metrics) in the limit $r=r_{c}\rightarrow\infty$. Crucially, it enables us to establish a bridge between the conventional Iyer–Wald formalism and the Euclidean background subtraction method  [^3] An alternative background metric choice with $\lambda(r_{c})=1$ exists; though useful for analyzing the Smarr relation, it is irrelevant to the present discussion.. For this horizonless background spacetime, we have $0=\int_{S_{\infty}}\bm{k}_{\xi_{t}}$, leading to
@@
For the Schwarzschild-AdS black hole example, one can insert the expressions of $\bm{Q}_{\xi_{t}}$ and $\bm{B}$ into Eq. (26) and verify that it indeed gives the expected mass $M=m$. This expression can also be rewritten in the ADM or Brown–York form, but that lies beyond our present scope; see Iyer:1994ys ; Harlow:2019yfa ; Guo:2024oey for details. Besides, the term $T\delta S$ in Eq.(25) appears to be non-integrable at first glance; however, it is indeed integrable, as the simplest Schwarzschild-AdS spacetime depends on only a single parameter. This fact becomes transparent when we express it in the form $T(M)\delta S(M)$.
@@
For asymptotically AdS black holes in Einstein gravity and higher-derivative gravity theories, Ref. Guo:2025ohn shows that the boundary term always vanishes after subtraction, following a careful examination of the asymptotic behaviors of such terms. This means that when using Eqs. (26) and (28) to compute the mass and Euclidean action, one may safely ignore $\bm{B}_{E}$. We remark that this property holds for pure gravity in AdS spacetime. Once we go beyond this regime, such as in matter-coupled theories, we should carefully re-examine possible boundary contributions.
@@ III Applying the subtraction method to matter–coupled gravity
In last section, we have shown that for pure gravity, the background subtraction method directly follows from the Iyer–Wald formalism and yields correct thermodynamics. The reasoning is quite general, so there is no obvious reason why this conclusion should fail when matter fields are present. For a diffeomorphism-invariant Lagrangian, inclusion of matter only modifies the explicit expressions of $\bm{Q}$ and $\bm{\Theta}$ by adding matter-related contributions, and none of the general steps in Sec. II are altered.
@@ III.1 Higher-derivative gravity with non-minimally coupled Maxwell field
$\bm{k}_{\xi}=\delta\mathbf{Q}_{\xi}-\xi\cdot\bm{\Theta}[\delta\phi],$ (38)
@@
We now take the timelike Killing vector $\xi_{t}=\frac{\partial}{\partial t}$. The identity $\int_{S_{h}}\bm{k}_{\xi_{t}}=\int_{S_{\infty}}\bm{k}_{\xi_{t}}$ produces
@@
Next, we introduce the reference background spacetime. For the metric in Eq. (34), we set $r_{h}=q=0$ and apply a redshift factor to ensure the black hole spacetime and background behave consistently as $r=r_{c}\to\infty$. For this horizonless background, the identity $\int_{S_{\infty}}\bm{k}_{\xi_{t}}=0$ holds, leading to:
@@ III.3.1 Non-integrable variations from matter field parameters
In matter-coupled gravity, variations of matter field parameters contribute to $\bm{k}_{\xi_{t}}=\delta\mathbf{Q}_{\xi_{t}}-\xi_{t}\cdot\bm{\Theta}[\delta\phi]$, which can lead to non-integrable variations, i.e., expressions that cannot be written as the differential of a single quantity.
For gauge fields, this issue is often resolved by a gauge choice: for example, if the gauge field $A_{\mu}$ is required to vanish at the horizon, one finds $\int_{\infty}\bm{k}_{\xi_{t}}=\delta M-\Phi\delta Q$, which is a non-integrable variation. But redefining the gauge can shift the $\Phi\delta Q$ term from the asymptotic boundary to the horizon, neatly yielding $\int_{\infty}\bm{k}_{\xi_{t}}=\delta M$.
@@
$\int_{\infty}\bm{k}_{\xi_{t}}\sim-2\delta f_{v}+\phi_{s}\delta\phi_{v}+2\phi_{v}\delta\phi_{s},$ (71)
@@
This non-integrability does not indicate a failure of the Iyer–Wald formalism and the subtraction method. Actually, they still yield the correct thermodynamic relations among variations of different quantities involving $\delta f_{v}$, $\delta\phi_{v}$, and $\delta\phi_{s}$. The main problem is that ambiguities arise when identifying $\delta M$ from $\int_{\infty}\bm{k}_{\xi_{t}}$: neither the Iyer–Wald formalism nor the background subtraction method provides a clear criterion for classifying these terms. Accordingly, one must resort to an alternative approach. In this regard, Ref. Li:2020spf resolves the issue using holographic renormalization, properly rewriting the integral (71) as
$\int_{\infty}\bm{k}_{\xi_{t}}\sim\delta\mathcal{E}+\langle O\rangle d\phi_{s},$ (74)
@@ III.3.2 Discrepancy between black hole entropy and Wald entropy
The black hole entropy $S$ is defined via $\int_{S_{h}}\bm{k}_{\xi}=T\delta S$, and the Wald entropy $S_{W}$ is defined via $\int_{S_{h}}\mathbf{Q}_{\xi}=TS_{W}$, where $\xi$ has to be taken as the horizon Killing vector of the spacetime under study. These two quantities coincide with each other, differing only in extreme cases where matter fields diverge at the horizon An:2024fzf ; Feng:2015oea .
Concretely, the equality between $S$ and $S_{W}$ relies on the horizon Killing vector vanishing at the bifurcation surface, which eliminates the $\xi\cdot\bm{\Theta}[\delta\phi]$ term in $\bm{k}_{\xi}$. However, if a matter field diverges at $r=r_{h}$, it can counteract this effect and yield a non-trivial contribution to $\int_{S_{h}}\bm{k}_{\xi}$, causing the inequality $S\neq S_{W}$. When this situation arises, it brings about difficulties in consistently interpreting black hole thermodynamics. The awkward problem of $S_{W}\neq S$ (within the Iyer–Wald formalism) translates to $I_{E}=\beta(M-TS_{W})\neq\beta(M-TS)$ (within the Euclidean action method). In such cases, $I_{E}$ cannot be legitimately related to the system’s free energy, and we thus cannot extract the various thermodynamic quantities of the black hole from $I_{E}$.
@@ IV Concluding remarks
We note here that, although we focus on AdS spacetimes in Sec. II, all of our analysis also applies to asymptotically-flat spacetimes. In such scenarios, the boundary term $\int_{\partial\mathcal{M}}^{(\mathrm{Reg.})}\bm{B}_{E}$ yields the dominant contribution to the Euclidean action $I_{E}$. For example, for a Schwarzschild black hole in Einstein gravity, the bulk action vanishes identically, and the Euclidean action takes the form $\int_{\partial\mathcal{M}}^{(\mathrm{Reg.})}\bm{B}_{E}\sim\int_{\partial M}(K-K_{0})\sqrt{|h|}d^{3}x$. We emphasize that the last term involves $K_{0}\sqrt{|h|}$ rather than $K_{0}\sqrt{|h_{0}|}$, as the background spacetime and the black hole spacetime share the same asymptotic geometry at infinity, which is the key property guaranteed by the redshift factor.
@@ Acknowledgments
YX would like to thank Hong-Bao Zhang, Xi-Yao Guo and Jie-Qiang Wu for the useful discussions. YX is also thankful to the Higgs Centre for Theoretical Physics at the University of Edinburgh for providing research facilities and hospitality during the visit. This work was supported in part by the National Natural Science Foundation of China with Grant No. 12475048, the Hebei Natural Science Foundation with Grant No. A2024201012, the Science Research Project of Hebei Education Department with Grant No. JCZX2026019, and the China Scholarship Council with Grant No. 202408130101.